OpenStudy (vera_ewing):

Joselyn is a manager at a sign painting company. She has three painters, Allen, Brianne, and Charles. Allen can complete a large project in 16 hours. Brianne can complete the same sized project in 18 hours. Charles is new, so no one knows how long it will take him. Joselyn assigns them all a large project to complete together. Explain to Joselyn how this project can tell her how long it would take Charles if he worked by himself. Use complete sentences.

OpenStudy (vera_ewing):

Can you explain to me how this project can see how long it would take Charles if he worked by himself? @saifoo.khan

OpenStudy (saifoo.khan):

We basically solve this by a general formula. \[\frac{1}{\text{time for each person on the job }}\] Then sum of these parts of the job add to one.

OpenStudy (vera_ewing):

But how do I explain? What would I say to Joselyn?

OpenStudy (saifoo.khan):

Like in this case, \[\frac{1}{\text{If all three work together}} = \frac{1}{\text{Allen's time}}+ \frac{1}{\text{Brianne's time}}+\frac{1}{\text{Charles's time}}\]

OpenStudy (vera_ewing):

Like in what case? Nothing is showing up!

OpenStudy (vera_ewing):

Hmm ok...so how do I explain it? What should I say for my answer? I have to use complete sentences

OpenStudy (saifoo.khan):

Start like how I did. Basically we fine the rate of each person for the job and then add them together.

OpenStudy (vera_ewing):

Then what? How do I explain it in complete sentences. That's what I need help with.

OpenStudy (saifoo.khan):

Read here. Sweetly written: http://www.purplemath.com/modules/workprob.htm

OpenStudy (saifoo.khan):

If you're still stuck, then come back here.